Consider the galvanic cell described by (N and M are metals):

N(s)|N2+(aq)||M+(aq)|M(s)

If Eocathode = 0.93 V and Eoanode = 1.814 V, and [N2+(aq)] = 0.568 M and [M+(aq)] = 0.299 M, what is Ecell, using the Nernst equation? ____ V

Provide your response to two digits after the decimal.
Answer #1

The given cell reaction is, N+2M* (aq)=N- (aq)+2M Anode: M M2 +2e Cathode: 2M*+2e +2M E° = E° cathode – E°anode =0.93 V-1.814